# SpecGen: Neural Spectral BRDF Generation via Spectral-Spatial Tri-plane Aggregation

### v1

@@ III-B Adaptive Feature Fusion (AFF)
$s={\mathcal{GAP}}\left({{\boldsymbol{f}_{h}}}\right)=\frac{1}{{H\times W}}\sum\limits_{i=1}^{H}{\sum\limits_{j=1}^{W}{\boldsymbol{f}_{h}(i,j)}}.$ (6)
To enrich feature representations, inspired by bottleneck structures [38, 39], we first expand $s$ into a descriptor $\boldsymbol{z}\in\mathbb{R}^{d}$ ($d=32$) using a $1\times 1$ convolution followed by a ReLU activation. We then apply six parallel, branch-specific convolutions, each with kernel $\boldsymbol{\beta}_{i}$, to reduce the channel dimensionality, enhance inter-channel dependencies, and extract multi-branch feature representations. The process for the $i$-th branch is as follows:
@@
The input spatial and spectral features are re-weighted by their corresponding adaptive scores and then aggregated to form the final fused output $\boldsymbol{f}_{u}\in\mathbb{R}^{C}$, which effectively captures the relative importance of different features during fusion:
@@
where $\boldsymbol{\delta}$ are the MLP parameters.
With SSTA feature representation, AFF-based fusion, and this MLP mapping, we predict the spectral reflectance $r$ at each coordinate. Repeating this process over all directions and wavelengths, we reconstruct the full spectral BRDF, enabling flexible spectral image rendering under diverse illumination and geometry.
III-C RGB-Spectral Joint Training Strategy
@@
$\mathcal{L}_{RGB}=\frac{1}{N}\sum_{i=1}^{N}(\hat{r}^{\prime}-g^{\prime})^{2},$ (16)
@@
To train SpecGen, we employ the loss functions $\mathcal{L}_{RGB}$ and $\mathcal{L}_{spec}$ for RGB and spectral BRDF data, respectively. This joint training strategy allows us to leverage large-scale RGB BRDF data, significantly enhancing the model’s generalization capability.
@@
Figure 4: Qualitative comparisons of reconstructed images under environmental illumination are shown alongside results from state-of-the-art hyperspectral image reconstruction methods, including HDNet [40] and MST++ [34]. Our method demonstrates superior spectral accuracy and robustness across different wavelength bands and lighting conditions, yielding reconstructed images that closely match the GT and outperform existing approaches in visual fidelity.
IV Experiment
IV-A Dataset and Evaluation Metrics
We employ two datasets in our experiments: RGL dataset [6] and MERL dataset [14]. RGL [6] dataset comprises 51 real-world isotropic materials and provides RGB and spectral BRDFs. MERL [14] dataset includes 100 real-world materials. Each material consists of a set of reflectance measurements, parameterized as spherical angles $(\theta_{h},\theta_{d},\varphi_{d})$ derived from the Rusinkiewicz parameterization [35] using the $\boldsymbol{h}$ and $\boldsymbol{d}$ vectors. The resolution for each wavelength channel is $90\times 90\times 180$, with a total of 1,458,000 reflectance measurements across 195 wavelength channels. Instead of directly comparing tabulated BRDF values, we render the predicted spectral BRDFs on a reference sphere under diverse illumination conditions and compute PSNR [41] and SSIM [41] against images rendered with the GT BRDFs, ensuring that the scores reflect perceptual fidelity in practical use cases.
[…]
Figure 5: Rendered images of the same material across additional wavelengths (400–1000 nm), using the same input as in Fig. 1. Comparison with the GT further demonstrates the accuracy of our method.
IV-B Implementation Details
We implement all experiments using PyTorch with the Adam optimizer [42], setting $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is initialized at $1\times 10^{-4}$, batch size is 65,536 (sampled points from multiple materials), and the learning rate decays by half every four epochs for stable convergence. Our model trains for 20 epochs on an RTX 4090 GPU (approx. 20 hours). We use 36 materials from the RGL isotropic dataset [6] for training and 12 for testing. For both, we sample 5,120,000 points per material (about 1/10 of total) and apply $\mu$-law compression to reduce dynamic range. Spectral and spatial triplane dimensions are initialized as [90, 90, 180, 39], with 39 for wavelength channel.
IV-C Evaluation on Spectral BRDF Generation
Baselines
Our method effectively reconstructs spectral BRDFs of various materials from just a single input image, enabling detailed rendering of spectral images at different wavelengths for diverse purposes. In contrast, hyperspectral image reconstruction methods directly recover spectral images from input images with a different focus. Although their objectives differ, both produce spectral representations evaluable for fidelity and overall accuracy. Thus, we assess our model’s performance by comparing the quality of rendered spectral images using well-established metrics like PSNR [41] and SSIM [41] under controlled experimental conditions. We compare with state-of-the-art hyperspectral methods, including HD-Net [40], for high-resolution recovery, and MST++ [34], using a mask-guided transformer for spectral reconstruction.
Table II: Quantitative comparisons on spectral rendering between our method and existing hyperspectral image reconstruction methods under environmental lights (top) and distant lights (middle).
Material HD-Net [40] MST++ [34] Ours
@@ IV-C Evaluation on Spectral BRDF Generation
Qualitative and Quantitative Assessment
As shown in Table II, for three representative materials, we input images of spheres rendered under both environmental and distant lights and use our method alongside baseline models to render spectral images at nine uniformly selected wavelengths for quantitative comparison. The results demonstrate that our method surpasses the baselines in multispectral image reconstruction under varying lighting conditions, achieving an average PSNR improvement of over 8 dB. Furthermore, as shown in Fig. 3, our method consistently produces spectral reconstructions that are much closer to the GT than HSI reconstruction methods, not only under parallel lighting but also across a wide range of environmental illumination conditions and spectral bands. Additional qualitative results in Fig. 4 further demonstrate the superior generalization and robustness of our approach in challenging environment light scenarios, with accurate rendering across the visible and near-infrared spectra.
Table III: Ablation studies on the robustness of our method under different environmental lights. Gray cells indicate environmental lights present in training datasets. Our method shows generalization even in environmental light that is not covered by the training dataset.
@@
[…]
Figure 6: Spectral BRDF results for “acrylic-felt-yellow” generated from inputs rendered under diverse unseen environmental lights (Outdoor, Indoor1), with outputs rendered under distant light. The left side displays the environmental lights used for rendering the input images. Our method demonstrates robust performance across varying environmental lighting conditions, closely matching the GT across wavelengths.
[…]
In addition, while existing HSI methods are typically constrained to the 400-700 nm wavelength range and can only generate hyperspectral images of the same shape, our method enables spectral rendering over a broader wavelength spectrum and produces spectral BRDFs for multispectral rendering of different objects. Fig. 5 presents additional results of spectral renderings for various materials at representative wavelengths spanning the visible and near-infrared regions, demonstrating the generalization ability and high fidelity of our approach across a wider spectral range. Moreover, as shown in Fig. 3, after deriving the spectral BRDF from the input sphere image, our method can render the multispectral image of the bunny shape, achieving better reconstruction results over a broad wavelength range.
Table IV: Evaluation on the effectiveness of boosting spectral BRDF generation via RGB BRDF data on 12 test spectral BRDF data. We calculate PSNR and SSIM of spectral rendering of a sphere using GT and generated spectral BRDFs under distant light.
@@
Robustness Assessment
In real-world scenarios, materials are often illuminated under varying environmental lighting conditions. A robust reconstruction method should maintain high recovery fidelity regardless of these variations. As shown in Table LABEL:tab:_environmental_light, we evaluate our model using three different environmental lighting setups: Outdoor Env. map, Indoor Env. map1, and Indoor Env. map2. Notably, Indoor Env. map2 was included in the training datasets, while the other two were unseen during training. We rendered reconstructed images at 20 different wavelengths, computing PSNR and SSIM as the average across these wavelengths to assess reconstruction quality. The result that our method still performs well under previously unseen illumination shows that our model exhibits strong robustness to different environmental lighting. Additionally, as illustrated in Fig. 6, qualitative visualizations confirm this robustness in the same material, with our method producing realistic and highly accurate renderings across different wavelengths, closely matching the GT.
Table V: Ablation studies on triplane feature fusion modules, including Hadamard product used in K-Planes [15] and our proposed AFF module.
@@
IV-D Ablation study
Training strategy
To evaluate the effectiveness of our RGB-spectral joint training strategy in addressing the data scarcity problem, we conduct ablation studies comparing our method with and without the inclusion of RGB BRDF data during training. Specifically, we define two training sets: (1) RGL-S, which contains 36 spectral BRDFs from the RGL dataset [6]; and (2) RGL-C-S+MERL, which comprises RGL-S, 51 additional RGB BRDFs (RGL-C) from the RGL dataset [6], and an extra 100 RGB BRDFs from the MERL dataset [14]. We train SpecGen on these two datasets and evaluate the model on 12 unseen spectral BRDFs. As shown in Table LABEL:tab:_RGL-C-S+MERL, integrating RGB data into the training set significantly improves spectral reconstruction accuracy, yielding an average PSNR improvement of nearly 0.9 dB compared to using only spectral data.
Furthermore, we provide a qualitative comparison in Fig. 7, showcasing results across seven different wavelengths for two materials. The qualitative results demonstrate that incorporating RGB data into training enables our proposed model to predict reflectance more accurately across various wavelengths, producing rendered images that closely match the GT. This experiment result efficiently highlights the effectiveness of leveraging RGB-BRDF data to enhance spectral BRDF reconstruction, further demonstrating the rationalization of our training strategy.
@@ IV-D Ablation study
Previous studies [15] have highlighted the effectiveness of the Hadamard product in triplane feature fusion. To evaluate whether our proposed Adaptive Feature Fusion (AFF) module offers a more efficient alternative, we conduct a correlation ablation study in Table LABEL:table:_aff. The ablation results demonstrate a significant advantage of our AFF in multiplanar feature fusion, achieving an average PSNR improvement of 4.3 dB across two test datasets compared to the Hadamard product strategy, showing the rationality and strength of the adaptive selection fusion mechanism in our AFF module.
V Conclusion
We propose SpecGen, the first framework for generating spectral BRDFs from a single RGB image of a sphere, enabling photorealistic spectral rendering under arbitrary illuminations and geometries. Our method addresses the fundamental challenge of spectral BRDF data scarcity by using abundant RGB BRDF data to enhance spectral BRDF generation, which is achieved by our Spectral-Spatial Tri-plane Aggregation (SSTA) network that decouples reflectance modeling into spectral and spatial domains. Furthermore, our Adaptive Feature Fusion (AFF) module effectively integrates multi-plane features, overcoming the limitations of fixed fusion operators. Extensive experiments show the performance of our SpecGen on publicly available spectral BRDF datasets and downstream spectral rendering tasks.
VI Limitations and Future Work
Although the proposed SpecGen framework delivers encouraging results, there are a few aspects that could benefit from additional investigation.
Limitations
First, since the sampling patterns used during training do not fully match the illumination distributions at test time, renderings under complex environment maps occasionally exhibit mild artifacts, as shown in Fig. 4. Existing neural sampling techniques developed for RGB BRDFs provide valuable insights, yet the higher dimensionality of spectral BRDFs means that these RGB-oriented strategies do not always carry over directly. Second, even with the aid of a total-variation term that promotes smoothness, faint noise can still appear on highly continuous surfaces (e.g., spheres). We attribute this mainly to the present tri-plane formulation, which may be less expressive for extremely smooth functions. Third, the current prototype is tailored to a single RGB photograph; this design choice simplifies network conditioning but also restricts immediate applicability to scenes that feature more complex geometry or spatially varying materials.
Future Work
We plan to extend SpecGen toward scene-level spectral BRDF estimation so that spatially varying materials in real-world environments can be reconstructed from casual multi-view imagery. Another promising direction is the development of neural sampling schemes specifically tuned to the spectral domain, which may better accommodate its higher dimensional characteristics. We are also interested in incorporating adaptive wavelength discretization strategies to cover broader spectral ranges and to model anisotropic reflectance effects. Finally, we will explore additional regularization techniques, more flexible multi-plane decompositions, and lightweight model variants to reduce memory footprint and inference time while further suppressing residual noise and improving output smoothness.

### v2

@@ 3.2 Adaptive Feature Fusion (AFF)
$s={\mathrm{GAP}}\left({{\boldsymbol{f}_{h}}}\right)=\frac{1}{{H\times W}}\sum\limits_{i=1}^{H}{\sum\limits_{j=1}^{W}{\boldsymbol{f}_{h}(i,j)}}.$ (6)
To enrich feature representations, inspired by bottleneck structures [hu2018squeeze, li2024efficient], we first expand $s$ into a descriptor $\boldsymbol{z}\in\mathbb{R}^{d}$ ($d=32$) using a $1\times 1$ convolution followed by a ReLU activation. We then apply six parallel, branch-specific convolutions, each with kernel $\boldsymbol{\beta}_{i}$, to reduce the channel dimensionality, enhance inter-channel dependencies, and extract multi-branch feature representations. The process for the $i$-th branch is as follows:
@@
Input spatial and spectral features are re-weighted by their corresponding adaptive scores and then aggregated to form the final fused output $\boldsymbol{f}_{u}\in\mathbb{R}^{C}$, which effectively captures the relative importance of different features during fusion:
@@
where $\boldsymbol{\delta}$ are the MLP parameters. With SSTA feature representation, AFF-based fusion, and this MLP mapping, we predict the spectral reflectance $r$ at each coordinate. Repeating this process over all directions and wavelengths, we reconstruct the full spectral BRDF, enabling flexible spectral image rendering under diverse illumination and geometry.
3.3 RGB-Spectral Joint Training Strategy
@@
$L_{\mathrm{RGB}}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{r}^{\prime}_{i}-g^{\prime}_{i}\right)^{2}$ (16)
@@
To train SpecGen, we employ the loss functions $\mathcal{L}_{RGB}$ and $\mathcal{L}_{spec}$ for RGB and spectral BRDF data, respectively. This joint training strategy allows us to leverage large-scale RGB BRDF data, enhancing the model’s generalization.
4 Experiment
4.1 Dataset and Evaluation Metrics
We evaluate our method on two public BRDF datasets: the RGL dataset [dupuy2018adaptive] and the MERL dataset [matusik2003data]. RGL contains 51 real-world isotropic materials with both RGB and spectral BRDF measurements. MERL provides BRDFs for 100 real-world materials, parameterized in the Rusinkiewicz domain $(\theta_{h},\theta_{d},\varphi_{d})$ [rusinkiewicz1998new] using the $\mathbf{h}$ and $\mathbf{d}$ vectors. Each wavelength channel has a resolution of $90\times 90\times 180$, yielding 1,458,000 measurements across 195 wavelengths. To enable meaningful perceptual comparison, we render predicted spectral BRDFs on a reference sphere under diverse illuminations and compute PSNR and SSIM [wang2002universal] against renderings generated from ground-truth BRDFs.
4.2 Implementation Details
We implement all experiments using PyTorch with the Adam optimizer [kingma2014adam], setting $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is initialized at $1\times 10^{-4}$, batch size is 65,536 (sampled points from multiple materials), and the learning rate decays by half every four epochs for stable convergence. Our model trains for 20 epochs on an RTX 4090 GPU (approx. 20 hours). We use 36 materials from the RGL isotropic dataset [dupuy2018adaptive] for training and 12 for testing. For both, we sample 5,120,000 points per material (about 1/10 of total) and apply $\mu$-law compression to reduce dynamic range. Spectral and spatial triplane dimensions are initialized as [90, 90, 180, 39], with 39 for wavelength channel.
4.3 Evaluation on Spectral BRDF Generation
Baselines
Our method reconstructs spectral BRDFs of diverse materials from a single input image, enabling detailed spectral rendering across wavelengths. In contrast, hyperspectral image reconstruction methods aim to recover spectral images directly, with a different modeling focus. Despite this distinction, both generate spectral representations that can be evaluated for fidelity. Therefore, we assess the quality of rendered spectral images using PSNR and SSIM [wang2002universal] under controlled conditions. We further compare our approach with state-of-the-art hyperspectral methods, including HD-Net [hu2022hdnet] for high-resolution recovery and MST++ [cai2022mask], a mask-guided transformer for spectral reconstruction.
@@ 4.3 Evaluation on Spectral BRDF Generation
Qualitative and Quantitative Assessment
As shown in Table 2, for three representative materials we compare our method with baselines using sphere images rendered under both environmental and distant lights, generating spectral images at nine uniformly spaced wavelengths. Our method consistently outperforms in multispectral reconstruction across diverse lighting conditions, with an average PSNR gain of over 8 dB. As shown in Fig. 3, it yields reconstructions significantly closer to the ground truth than HSI methods under both parallel and varied environmental lighting. Supplementary material provides additional qualitative results that further confirm its superior generalization ability and robustness in complex lighting scenarios, with accurate rendering across the visible and near-infrared spectra.
Unlike existing HSI methods, typically limited to the 400–700 nm visible spectrum and constrained to fixed-shape outputs, our method supports a significantly broader spectral range covering both visible and near-infrared bands, and generates high-fidelity spectral BRDFs that enable realistic multispectral rendering of objects with diverse shapes and materials. Supplementary material provides results for various materials at representative visible and near-infrared wavelengths, demonstrating strong generalization and fidelity. As shown in Fig. 3, using the spectral BRDF from a sphere image, our method can render a bunny model with superior reconstruction across a wide wavelength range.
Robustness Assessment
Materials in real-world scenes face diverse environmental lighting. A robust reconstruction should preserve fidelity under such variations. As shown in Table LABEL:tab:_environmental_light, we evaluate our model under three lighting setups: Outdoor Env. map, Indoor Env. map1, and Indoor Env. map2. Only the last was seen in training. We reconstruct images at 20 wavelengths and report the average PSNR and SSIM to assess quality. Strong performance on unseen lighting shows the model’s robustness. Supplementary material provides qualitative results showing realistic, accurate renderings across wavelengths that closely match the GT.
Table 2: Quantitative comparisons on spectral rendering between our method and existing hyperspectral image reconstruction methods under environmental lights (top) and distant lights (middle).
Material HD-Net [hu2022hdnet] MST++ [cai2022mask] Ours
@@
Table 3: Ablations on the robustness of our method under different environmental lights. Gray cells indicate environmental lights present in training datasets. Our method shows generalization even in environmental light that is not covered by the training dataset.
@@
4.4 Ablation study
Training strategy
To assess our RGB–spectral joint training strategy for mitigating data scarcity, we compare models trained with and without RGB BRDF data. We use two training sets: (1) RGL-S, containing 36 spectral BRDFs from RGL [dupuy2018adaptive]; and (2) RGL-C-S+MERL, which adds 51 RGB BRDFs from RGL and 100 from MERL [matusik2003data]. SpecGen is trained on each and evaluated on 12 unseen spectral BRDFs. As shown in Table LABEL:tab:_RGL-C-S+MERL, adding RGB data improves spectral reconstruction by nearly 0.9 dB PSNR on average. Figure 4 provides qualitative results for two representative materials at seven distinct wavelengths, showing that RGB–spectral joint training yields more accurate reflectance predictions and produces rendered images that are perceptually and quantitatively closer to the GT. These results highlight the effectiveness of leveraging abundant RGB BRDF data as complementary supervision to enhance the fidelity and consistency of spectral BRDF reconstruction across the spectrum.
Table 4: Evaluation on the effectiveness of boosting spectral BRDF generation via RGB BRDF data on 12 test spectral BRDF data. We calculate PSNR and SSIM of spectral rendering of a sphere using GT and generated spectral BRDFs under distant light.
@@ 4.4 Ablation study
Table 5: Ablations on triplane feature fusion modules, including Hadamard product used in K-Planes [fridovich2023k] and our AFF module.
@@
Previous studies [fridovich2023k] have highlighted the effectiveness of the Hadamard product in triplane feature fusion. To evaluate whether our proposed Adaptive Feature Fusion (AFF) module offers a more efficient alternative, we conduct a correlation ablation study in Table LABEL:table:_aff. The ablation results demonstrate a significant advantage of our AFF in multiplanar feature fusion, achieving an average PSNR improvement of 4.3 dB across two test datasets compared to the Hadamard product strategy, showing the rationality and strength of the adaptive selection fusion mechanism in our AFF module.
5 Limitations and Future Work
Despite its strong empirical performance, SpecGen exhibits several limitations. First, the current framework assumes spherical input geometry. Relaxing this assumption to accommodate arbitrary shapes may lead to measurable accuracy degradation. Second, materials with pronounced specular behavior are highly sensitive to illumination variations, complicating reliable recovery of their spectral BRDFs. Third, naively extending RGB-based BRDF generation methods to the spectral domain substantially increases model complexity, undermining the fairness and interpretability of cross-method comparisons. Finally, while SpecGen reconstructs full BRDFs from sparse samples, the limited size of existing spectral BRDF datasets constrains its upper-bound performance. We plan to investigate the performance of our method once more spectral BRDFs are available, and compare with more baselines in our future work.
6 Conclusion
We propose SpecGen, the first framework that generates spectral BRDFs from a single RGB sphere image, enabling photorealistic spectral rendering under arbitrary illuminations and geometries. To address the scarcity of spectral BRDF data, SpecGen leverages abundant RGB BRDFs through a SSTA that decouples reflectance modeling into spectral and spatial domains. Moreover, our AFF module integrates multi-plane features more effectively than fixed fusion operators. Extensive experiments on public spectral BRDF datasets and downstream spectral rendering tasks demonstrate the strong performance of our SpecGen.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62472044, U24B20155), Hebei Natural Science Foundation Project No. 242Q0101Z, Beijing-Tianjin-Hebei Basic Research Funding Program No. F2024502017, and the Youth Innovative Research Team of BUPT.